OpenStudy (isabel☺):

how to integrate:

OpenStudy (isabel☺):

\[\pi \int\limits y \sin^2(y) dy\]

OpenStudy (anonymous):

You'd really want to use the double angular formula here for the sin^2(y) term

OpenStudy (isabel☺):

1/2 -1/2 cos 2y?

OpenStudy (anonymous):

yes that's the one.

OpenStudy (isabel☺):

\[\pi \int\limits y (1/2 -1/2 \cos 2y) dy\]

OpenStudy (isabel☺):

\[1/2 \pi \int\limits (y - ycos2y) dy\] i dont know that part: \[\int\limits y \cos 2y dy\]

OpenStudy (chillout):

Try integrating by parts!

OpenStudy (isabel☺):

\[1/4 \pi y^2- 1/4 \pi y siny - 1/8 \pi \cos 2y\] ?

OpenStudy (chillout):

That's right.

OpenStudy (chillout):

Oh, and it's sin(2y)!